Use the normal distribution of IQ​ scores, which has a mean of 125 and a standard deviation of 11,and the following table with the standard scores and percentiles for a normal distribution to find the indicated quantity.

Percentage of scores greater than 152.5 is ___ %

Standard score Percent
-3 0.13
-2.5 0.62
-2 2.28
-1.5 6.68
-1 15.87
-0.9 18.41
-0.5 30.85
-0.1 46.02
0 50
0.1 53.98
0.5 69.15
0.9 81.59
1 84.13
1.5 93.32
2 97.72
2.5 99.38
3 99.87
3.5 99.98
Answer #1

Solution :

Given that,

Mean = 125

SD = 11

P(x > 152.5) = 1 - P(x < 152.5)

= 1 - P((x - mean) / sd < (152.5 - 125) / 11))

= 1 - P(z < 2.5)

= 1 - 0.9938

= 0.0062

Answer = 0.62%